OpenStudy (coexist2):

A slide 4.4 m long makes an angle of 33 degrees with the ground. How high is the top of the slide above the ground? A. 2.53m B. 2.4m C. 3.69m D. 2.86m

OpenStudy (mecharv):

Sin(33)=Height/4.4

OpenStudy (mecharv):

Can you find height from the above relation? @CoExist2

OpenStudy (coexist2):

0.5446390

OpenStudy (coexist2):

?

OpenStudy (mecharv):

Multiply it with 4.4

OpenStudy (mecharv):

Height= sin(33 degrees)*4.4

OpenStudy (abdullahm):

 Hint: Use sin \( \Large\sf sin \theta = \frac{opposite~side}{hypotenuse}\)

OpenStudy (abdullahm):

Drawing the picture is helpful and let's you solve the problem faster :)

OpenStudy (coexist2):

thanks.

OpenStudy (abdullahm):

It was my pleasure!

OpenStudy (jesselacompte):

c